Harrington Corporation produces three products, A, B, and C. Pertinent information on these products is as follows:

Product Selling Price per Unit Variable Cost per Unit Fixed Cost
per Unit
DL Hours
per Unit
A (Anchor bolts) $ 8.00 $ 2.00 $ 3.00 1
B (Bearings) $ 3.60 $ 2.60 $ 4.00 2
C (Castings) $ 8.00 $ 4.00 $ 1.00 4

There are 160 direct labor hours available. Machine-hour capacity allows 100 anchor bolts, only; 40 bearings, only; 60 casters, only; or any combination of the three that does not exceed the capacity.

The machine-hour constraint for Harrington's linear programming model is:

Multiple Choice

  • The same as the labour-hour constraint.

  • 1A + 2.50B + 1.67C ≤ 100.

  • Indeterminable from the data given.

  • A + B + C ≤ 200.

  • 100A + 40B + 60C ≤ 20,000.
Answer #1
Ratio of bearings = 100/40 = 2.50
Ratio of casters = 100/60 = 1.67
The machine-hour constraint for Harrington's linear programming model is:
1A + 2.50B + 1.67C ≤ 100.
Option B is correct